Find the initial speed which a projectile must be subjected to give it a maximum horizontal range of 490m. Assume the acceleration due to gravity as g=10m/s


1
Expert's answer
2021-05-03T10:28:16-0400
"R=\frac{v^2}{g}\\490=\frac{v^2}{10}\\v=70\frac{m}{s}"